### v1

# Galactokinetics

Abstract
Galactic disks lie at the heart of many of the most pressing astrophysical puzzles. There are sophisticated kinetic theories that describe some aspects of galaxy disk dynamics, but extracting quantitative predictions from those theories has proven very difficult, meaning they have shed little light on observations/simulations of galaxies. Here, we begin to address this issue by developing a tractable theory describing fluctuations and transport in thin galactic disks. Our main conceptual advance is to split potential fluctuations into asymptotic wavelength regimes relative to orbital guiding radius and epicyclic amplitude (similar to plasma gyrokinetics), and then to treat separately the dynamics in each regime. As an illustration, we apply our results to quasilinear theory, calculating the angular-momentum transport due to a transient spiral. At each stage we verify our formulae with numerical examples. Our approach should simplify many important calculations in galactic disk dynamics.
1 Introduction
Galactic disks occupy an especially prominent place in modern astrophysics, as the setting for many of our most profound puzzles. The nature of dark matter is still a mystery half a century after the discovery of flat rotation curves (Mo et al., 2010). Our understanding of galaxy formation in the early Universe is being upended by the observation of very massive early galaxies by JWST (Castellano et al., 2022; Wu et al., 2022). The assembly history and evolution of the Milky Way is still unknown although its fingerprints are surely embedded in the exquisite GAIA data (Deason & Belokurov, 2024). For these and other quandaries to be solved, we need an improved understanding of the dynamics and evolution of galactic disks.
The theorists’ picture of a galactic disk is a collection of stars on nearly circular, nearly coplanar orbits, which exchange energy and angular momentum with a bath of potential ‘fluctuations’. These fluctuations may comprise spiral transients, a central bar, dark matter substructure, interstellar medium (ISM) gas clouds, and so on. Given the importance of galaxy disk evolution it is highly desirable to develop kinetic theories that describe this energy and angular-momentum exchange accurately. Indeed, a great deal of sophisticated formalism has been developed for just this purpose, beginning several decades ago (Julian & Toomre, 1966; Kalnajs, 1971; Lynden-Bell & Kalnajs, 1972), long before observations of galaxy disks were of high enough quality to demand it, and long before the evolution of such disks could be simulated accurately on computers. For an overview, see Hamilton & Fouvry (2024).
And yet, despite this head start, there has been little progress in connecting abstract kinetic theory to either observation or numerical simulations. Milky Way astronomy is a prime example. GAIA has provided an unprecedented kinematic and chemical map of the Galaxy within a few kpc of the Sun, and several puzzles have thereby arisen regarding the formation of the vertical phase space ‘snail’ (Antoja et al., 2014), the remarkably efficient radial migration of disk stars (Frankel et al., 2022), and so on. Each of these puzzles has inspired various solutions, which normally consist of a plausible set of initial conditions and dynamical perturbations that are forward-modeled numerically, but in very few cases has this approach yielded a convincing answer that excludes all others: the parameter space of possibilities is simply too large for such models to land a knockout blow. But analytic (or semi-analytic) kinetic theory has not really helped either. With a few exceptions (e.g. Chiba & Schönrich 2021; Banik et al. 2022; Widrow 2023), kinetic theorists have been unable to write down tractable expressions — or even robust estimates or scaling laws — that can be meaningfully compared against observation. One could make similar remarks regarding the puzzles arising out of JWST, SDSS, etc. In short, the quality of modern observations demands a tractable, predictive kinetic theory of fluctuations and transport in galactic disks which is currently lacking.
The issue is that, although the formal kinetic equations describing disk evolution are (we think) largely known, extracting quantitative predictions from those equations is very difficult. For instance, one must almost always evaluate the $12d+$ 1-dimensional function
$\delta\phi_{\bm{n}}(\bm{J},t),$ (1)
where $d$ is the number of spatial dimensions ($d=2$ for a razor-thin disk, which is what we will concentrate on here). Really this is just the potential fluctuation $\delta\phi(\bm{r},t)$, except that it is (i) first transformed from position ($\bm{r}$) to angle-action space ($\bm{\theta},\bm{J}$), and (ii) then (Fourier) transformed from angle-space to wavenumber ($\bm{n}$) space. Furthermore, in many applications this function then undergoes (iii) a projection onto a set of biorthogonal basis functions, and then possibly (iv) yet another (Fourier or Laplace) transform from the time domain into the frequency domain, along with any necessary analytic continuation. After further manipulations there may even follow (v) an inverse Fourier or Laplace transform back to the time domain, likely requiring contour integration. This done, one typically must (vi) sum over all possible wavenumbers $\bm{n}$, of which there are in principle infinitely many. These devilish layers of additional complexity mean that in practice it is often computationally cheaper to simulate a system’s entire history directly than to evaluate an approximate ‘analytic’ expression designed to foretell its behavior over a single timestep. Such a theory can hardly be called predictive.
Luckily, things need not be so complicated. What we really care about is not the potential fluctuations themselves, but rather how they couple to orbits. In the case of galactic disks, there are two key orbital length scales — the guiding center radius $R_{\mathrm{g}}$ and the typical epicyclic amplitude $a$, with $a\ll R_{\mathrm{g}}$ in the common case of a ‘cool’ disk — and the potential fluctuations $\delta\phi(\bm{r},t)$ can be classified relative to these length scales. As we will show, in the resulting asymptotic regimes, one can derive expressions for $\delta\phi_{\bm{n}}(\bm{J},t)$ that are drastically simpler than in the general case referred to above, but also remarkably accurate in practice. Moreover, for some calculations the fluctuations from one asymptotic regime completely dominate the orbital couplings, so complicated kinetic equations can be greatly simplified. Our approach is reminiscent of the gyrokinetic theory of plasmas, in which the ‘mean-field’ orbits of particles consist of Larmor motion around magnetic field lines and force fluctuations are categorized relative to the Larmor radius. By analogy, we give our framework the name galactokinetics.
The purpose of this paper is to develop the galactokinetic approximation scheme and show how it renders tractable an important problem in the (quasilinear) kinetic theory of galactic disk dynamics. While some — perhaps even most — of the equations we develop below can be found scattered across the galaxy dynamics literature (Shu, 1970; Lynden-Bell & Kalnajs, 1972; Mark, 1974; Carlberg & Sellwood, 1985; Sygnet et al., 1988), the relevant approximation scheme has never been developed systematically before, nor does it appear to have been tested numerically. In a follow-up paper we will apply the formal techniques we develop here to the problem of the heating of the Milky Way disk.
Our plan for the rest of this work is as follows. First we introduce our basic model of a thin galactic disk (§2) and discuss the epicyclic approximation (§3). We then classify potential fluctuations according to their wavelengths, derive explicit expressions for $\delta\phi_{\bm{n}}(\bm{J},t)$ in each class, and validate the expressions numerically (§4). Having done this, we apply our asymptotic results to quasilinear kinetic theory (§5). We conclude in §6.
2 Model
We consider a razor-thin stellar disk made up of equal-mass stars (and possibly embedded in a rigid dark matter halo). We will describe the disk using polar coordinates $\bm{r}=(\varphi,R)$, and hence velocities using $\bm{v}=(v_{\varphi},v_{R})=(R\dot{\varphi},\dot{R})$. For simplicity, we will ignore any ‘vertical’ motion (motion normal to the disk plane).
To begin our construction, at a given time let the exact gravitational potential of the whole two-dimensional system (including external perturbations, if they are present) be $\phi(\bm{r})=\phi(\varphi,R)$. Then the exact Hamiltonian of a test star is
$h(\bm{r},\bm{v})=\frac{1}{2}\bm{v}^{2}+\phi(\bm{r}).$ (2)
The azimuthal average of $h$ is
$\widetilde{h}(R,\bm{v})\equiv\frac{1}{2}\bm{v}^{2}+\widetilde{\phi}(R),$ (3)
where for any function $q(\bm{r})$ we define
$\widetilde{q}(R)\equiv\frac{1}{2\pi}\int\mathrm{d}\varphi\,q(\bm{r}).$ (4)
We have not yet stipulated where the Hamiltonian $h$ comes from; it might be generated self-consistently by the stars, or externally imposed, or some combination of these.
We suppose that $\phi$ is reasonably close to axisymmetry, i.e., not far from $\widetilde{\phi}$, and we can construct angle-action variables $(\bm{\theta},\bm{J})$ such that the associated Hamiltonian $\widetilde{h}$ only depends on $\bm{J}$. These actions are $\bm{J}=(J_{\varphi},J_{R})$ where
$J_{\varphi}=Rv_{\varphi},$ (5)
is the angular momentum, and
$J_{R}=\frac{1}{\pi}\int_{R_{\mathrm{p}}}^{R_{\mathrm{a}}}\mathrm{d}R\,v_{R}(R),$ (6)
is the radial action. Here,
$v_{R}(R)\equiv\sqrt{2[E-\widetilde{\phi}(R)]-J_{\varphi}^{2}/R^{2}},$ (7)
is the radial velocity, and $R_{\mathrm{p/a}}$ are the peri/apocenter radii, i.e. the solutions to $v_{R}(R)=0$. The conjugate angles $\bm{\theta}=(\theta_{\varphi},\theta_{R})$ are then the azimuthal angle
$\theta_{\varphi}=\varphi+\int_{C}\mathrm{d}R^{\prime}\,\frac{\Omega_{\varphi}-J_{\varphi}/R^{\prime 2}}{v_{R}(R^{\prime})},$ (8)
and the radial angle
$\theta_{R}=\Omega_{R}\int_{C}\frac{\mathrm{d}R^{\prime}}{v_{R}(R^{\prime})},$ (9)
where
$\bm{\Omega}(\bm{J})=(\Omega_{\varphi},\Omega_{R})\equiv\frac{\partial\widetilde{h}}{\partial\bm{J}},$ (10)
are the orbital frequencies, and $C$ is the contour going from the pericenter $R_{\mathrm{p}}$ up to the current position $R(\theta_{R},\bm{J})$, along the radial oscillation. Thus $\theta_{R}$ increases by $2\pi$ in one radial period $T_{R}\equiv 2\pi/\Omega_{R},$ while $\theta_{\varphi}$ increases by $2\pi\Omega_{\varphi}/\Omega_{R}$.
2.1 Mean field and fluctuations
Having constructed our angle-action variables, we can express any function on phase space in terms of them via $\bm{r}=\bm{r}(\bm{\theta},\bm{J})$ and $\bm{v}=\bm{v}(\bm{\theta},\bm{J})$. In particular, we can return to our original exact Hamiltonian $h$ and split it into a $\bm{\theta}$-independent ‘mean-field’ part and a $\bm{\theta}$-dependent ‘fluctuation’:
$h(\bm{r}(\bm{\theta},\bm{J}),\bm{v}(\bm{\theta},\bm{J}))=h_{0}(\bm{J})+\delta h(\bm{\theta},\bm{J}),$ (11)
where the $\bm{\theta}$-average of some quantity $q(\bm{\theta},\bm{J})$ is
$q_{0}(\bm{J})\equiv\int\frac{\mathrm{d}\bm{\theta}}{(2\pi)^{2}}\,q(\bm{\theta},\bm{J}).$ (12)
Here, the ‘00’ subscript reminds us that $q_{0}$ is equivalent to the zeroth Fourier mode of $q$ when expanded in angles (see, e.g., equation (30) below). It is easy to show that the ‘mean-field Hamiltonian’ $h_{0}$ and the ‘axisymmetric Hamiltonian’ $\widetilde{h}$ are equivalent, and that
$\delta h(\bm{\theta},\bm{J})=\phi(\bm{r})-\widetilde{\phi}(R)\equiv\delta\phi(\bm{\theta},\bm{J}),$ (13)
i.e., the fluctuation part of the Hamiltonian is just the non-axisymmetric part of the potential. Hence, from now on we write
$h(\bm{r}(\bm{\theta},\bm{J}),\bm{v}(\bm{\theta},\bm{J}))=h_{0}(\bm{J})+\delta\phi(\bm{\theta},\bm{J}),$ (14)
and refer to $h_{0}$ and $\delta\phi$ as the mean-field Hamiltonian and the potential perturbation respectively.
Another key quantity we must introduce is the guiding center potential perturbation, which is defined as
$\delta\Phi(\theta_{\varphi},J_{\varphi})\equiv\delta\phi(\theta_{\varphi},\theta_{R},J_{\varphi},0).$ (15)
i.e., the potential fluctuation evaluated at the instantaneous location of the star’s guiding center at $J_{R}=0$. Note that this means $\delta\Phi$ is necessarily independent of $\theta_{R}$.
2.2 Distribution function
Let the exact distribution function (DF) of stars be $f(\bm{\theta},\bm{J},t)$, normalized such that
$\int\mathrm{d}\bm{\theta}\,\mathrm{d}\bm{J}f(\bm{\theta},\bm{J},t)=M,$ (16)
the total mass in stars. This DF satisfies the Klimontovich equation
$\frac{\partial f}{\partial t}+\frac{\partial f}{\partial\bm{\theta}}\cdot\frac{\partial h}{\partial\bm{J}}-\frac{\partial f}{\partial\bm{J}}\cdot\frac{\partial h}{\partial\bm{\theta}}=0.$ (17)
Without approximation we can always expand $f$ like we expanded $h$ in equation (14):
$f(\bm{\theta},\bm{J})=f_{0}(\bm{J})+\delta f(\bm{\theta},\bm{J}).$ (18)
(Recall that a subscript zero refers to an angle-average, viz. equation (12). Note also that here and henceforth we are suppressing the time argument of the fluctuations.). Plugging these expansions into (17) gives
$\frac{\partial\delta f}{\partial t}+\bm{\Omega}\cdot\frac{\partial\delta f}{\partial\bm{\theta}}-\frac{\partial f_{0}}{\partial\bm{J}}\cdot\frac{\partial\delta\phi}{\partial\bm{\theta}}$
$=-\frac{\partial}{\partial\bm{\theta}}\cdot\left(\delta f\frac{\partial\delta\phi}{\partial\bm{J}}\right)+\frac{\partial}{\partial\bm{J}}\cdot\left[\delta f\frac{\partial\delta\phi}{\partial\bm{\theta}}-\left(\delta f\frac{\partial\delta\phi}{\partial\bm{\theta}}\right)_{0}\,\right],$ (19)
and
$\frac{\partial f_{0}}{\partial t}=\frac{\partial}{\partial\bm{J}}\cdot\left(\delta f\frac{\partial\delta\phi}{\partial\bm{\theta}}\right)_{0}.$ (20)
We will frequently be interested in marginalized functions which do not depend on radial angle or radial action. To this end we introduce an overline notation to denote an integral (not an average) over $\theta_{R}$ and $J_{R}$:
$\overline{q}\equiv\int_{0}^{\infty}\mathrm{d}J_{R}\int_{0}^{2\pi}\mathrm{d}\theta_{R}\,q(\bm{\theta},\bm{J}).$ (21)
The most important function we wish to marginalize is the DF $f$. Integrating both sides of (18) with respect to $\theta_{R}$ and $J_{R}$, we have
$F(\theta_{\varphi},J_{\varphi})=F_{0}(J_{\varphi})+\delta F(\theta_{\varphi},J_{\varphi}),$ (22)
where
$F(\theta_{\varphi},J_{\varphi})$ $\equiv\overline{f(\bm{\theta},\bm{J})},$ (23)
$F_{0}(J_{\varphi})$ $\equiv\overline{f_{0}(\bm{J})}=2\pi\int_{0}^{\infty}\mathrm{d}J_{R}\,f_{0}(\bm{J}),$ (24)
$\delta F(\theta_{\varphi},J_{\varphi})$ $\equiv\overline{\delta f(\bm{\theta},\bm{J})}.$ (25)
By applying the same procedure to equations (2.2)–(20), we find that $\delta F$ and $F_{0}$ evolve according to
$\frac{\partial\delta F}{\partial t}+\overline{\left(\Omega_{\varphi}\frac{\partial\delta f}{\partial\theta_{\varphi}}\right)}$ $-\overline{\left(\frac{\partial\delta\phi}{\partial\theta_{\varphi}}\frac{\partial f_{0}}{\partial J_{\varphi}}\right)}$
$=-\frac{\partial}{\partial\theta_{\varphi}}\delta\mathcal{G}-\frac{\partial}{\partial J_{\varphi}}\delta\mathcal{F},$ (26)
and
$\frac{\partial F_{0}}{\partial t}=-\frac{\partial}{\partial J_{\varphi}}\mathcal{F}_{0},$ (27)
respectively, where
$\mathcal{F}(\theta_{\varphi},J_{\varphi})\equiv-\overline{\left(\delta f\frac{\partial\delta\phi}{\partial\theta_{\varphi}}\right)},$ (28)
$\mathcal{G}(\theta_{\varphi},J_{\varphi})\equiv\overline{\left(\delta f\frac{\partial\delta\phi}{\partial J_{\varphi}}\right)},$ (29)
and $\delta\mathcal{G}=\mathcal{G}-\mathcal{G}_{0}$, $\delta\mathcal{F}=\mathcal{F}-\mathcal{F}_{0}$ are the $\theta_{\varphi}$-dependent parts of $\mathcal{G}$ and $\mathcal{F}$. We emphasize that equations (26)–(29) are exact; we have made no approximations so far.
2.3 Fourier expansion
For explicit calculations it will be useful for us to introduce Fourier expansions of $\delta f$ and $\delta\phi$:
$\delta f(\bm{\theta},\bm{J})=$ $\sum_{{\bm{n}}}\delta f_{{\bm{n}}}(\bm{J})\,\mathrm{e}^{i{\bm{n}}\cdot\bm{\theta}},$ (30)
$\delta\phi(\bm{\theta},\bm{J})=$ $\sum_{{\bm{n}}}\delta\phi_{{\bm{n}}}(\bm{J})\,\mathrm{e}^{i{\bm{n}}\cdot\bm{\theta}},$ (31)
where $\bm{n}=(n_{\varphi},n_{R})$ are vectors of integers. In particular, various terms in equations (26)–(29) can then be expressed as Fourier series; some explicit expressions for these are given in Appendix A. The most important one is
$\mathcal{F}_{0}(J_{\varphi})=2\pi\sum_{\bm{n}}in_{\varphi}\int_{0}^{\infty}\mathrm{d}J_{R}\,\delta f_{\bm{n}}(\bm{J})\delta\phi^{*}_{\bm{n}}(\bm{J}).$ (32)
We will also benefit from developing the functions $\delta F$ and $\delta\Phi$ as Fourier series in the angle $\theta_{\varphi}$:
$\delta F(\theta_{\varphi},J_{\varphi})$ $=\sum_{n_{\varphi}}\mathrm{e}^{in_{\varphi}\theta_{\varphi}}\delta F_{n_{\varphi}}(J_{\varphi}),$ (33)
$\delta\Phi(\theta_{\varphi},J_{\varphi})$ $=\sum_{n_{\varphi}}\mathrm{e}^{in_{\varphi}\theta_{\varphi}}\delta\Phi_{n_{\varphi}}(J_{\varphi}).$ (34)
The coefficients in these expansions are related to those in (30)–(31) by
$\delta F_{n_{\varphi}}(J_{\varphi})=2\pi\int_{0}^{\infty}\mathrm{d}J_{R}\,\delta f_{(n_{\varphi},0)}(J_{\varphi},J_{R}),$ (35)
$\delta\Phi_{n_{\varphi}}(J_{\varphi})=\delta\phi_{(n_{\varphi},0)}(J_{\varphi},0).$ (36)
3 Epicyclic approximation
The results of §2 are exact, but to render the theory tractable we need to make some approximations. The key approximation is that the disk is cool (§3.1) which means that mean-field orbits are well-described by epicycles (§3.2).
3.1 Cool distribution function
We will assume throughout this study that our galactic disk is cool, by which we mean that at any given time, the underlying mean-field DF $f_{0}(\bm{J})$ is heavily biased towards near-circular orbits, i.e., those with $J_{R}\ll J_{\varphi}$. More precisely, we define the expectation value $\langle q\rangle$ of any quantity $q(\bm{\theta},\bm{J})$ as
$\langle q\rangle=\frac{1}{M}\int\mathrm{d}\bm{\theta}\mathrm{d}\bm{J}\,f(\bm{\theta},\bm{J})\,q(\bm{\theta},\bm{J}).$ (37)
Then the statement that the disk is cold is equivalent to the statement that, at all relevant angular momenta $J_{\varphi}$,
$\epsilon(J_{\varphi})\equiv\left(\frac{\langle J_{R}\rangle}{J_{\varphi}}\right)^{1/2}\ll 1.$ (38)
Usually we will drop the $J_{\varphi}$ dependence of the small quantity (38) and just refer to it as $\epsilon$.
How small is $\epsilon$ in reality? Recently, Binney & Vasiliev (2023) fit action-based DFs to various components of our Galaxy using GAIA data. They separated the Galactic disk into young, medium, old and thick components, each of which had a DF whose radial action dependence took the form $f_{0}\propto\exp(-J_{R}/\langle J_{R}\rangle)$, where $\langle J_{R}\rangle$ depends on $J_{\varphi}$. Using their best-fit models we deduce that at the Solar radius (assuming $J_{\varphi}=8\,\mathrm{kpc}\times 220\,\mathrm{km\,s}^{-1}$) these component disks have
$\epsilon(8\,\mathrm{kpc}\times 220\,\mathrm{km\,s}^{-1})=[0.11,0.13,0.13,0.17].$ (39)
Alternatively, evaluating $\epsilon$ somewhat closer to the Galactic center, we have
$\epsilon(4\,\mathrm{kpc}\times 220\,\mathrm{km\,s}^{-1})=[0.12,0.16,0.18,0.25],$ (40)
or somewhat further from the Galactic center, we have
$\epsilon(12\,\mathrm{kpc}\times 220\,\mathrm{km\,s}^{-1})=[0.10,0.11,0.11,0.14].$ (41)
Thus, at least for the regions of interest in the Milky Way, higher order ($\mathcal{O}(\epsilon^{2})$) corrections to the theory we are about to develop are at most a few percent.
3.2 Epicyclic orbits
When $\epsilon\ll 1$, most orbits are well described using the epicyclic approximation. The angle-action mapping for epicyclic orbits in a central potential $\widetilde{\phi}(R)$ is:
$\varphi$ $=\theta_{\varphi}+\gamma\frac{a_{R}}{R_{\mathrm{g}}}\sin\theta_{R},$ (42)
$R$ $=R_{\mathrm{g}}-{a_{R}}\cos\theta_{R},$ (43)
$J_{\varphi}$ $=R_{\mathrm{g}}^{2}\,\Omega_{\varphi},$ (44)
$J_{R}$ $=\frac{E_{R}}{\Omega_{R}},$ (45)
where $R_{\mathrm{g}}(J_{\varphi})$ is the guiding radius (the radius of a circular orbit of angular momentum $J_{\varphi}$), $a_{R}$ is the amplitude of the radial epicycle
$a_{R}(\bm{J})\equiv\frac{\sqrt{2E_{R}}}{\Omega_{R}},$ (46)
and $E_{R}$ is the radial energy
$E_{R}(\bm{J})\equiv\frac{1}{2}v_{R}^{2}+\frac{1}{2}\Omega_{R}^{2}(R-R_{\mathrm{g}})^{2}.$ (47)
The frequencies can be expressed in terms of $\widetilde{\phi}$ as
$\Omega_{\varphi}^{2}(\bm{J})$ $=\left(\frac{1}{R}\frac{\partial\widetilde{\phi}}{\partial R}\right)_{R=R_{\mathrm{g}}}\equiv\Omega^{2},$ (48)
$\Omega_{R}^{2}(\bm{J})$ $=\left(\frac{2\Omega_{\varphi}}{R}\frac{\partial(\Omega_{\varphi}R^{2})}{\partial R}\right)_{R=R_{\mathrm{g}}}\equiv\kappa^{2},$ (49)
and $\gamma$ is an order-unity dimensionless number
$\gamma\equiv\frac{2\Omega}{\kappa}.$ (50)
The azimuthal period is then approximately $T_{\varphi}=2\pi/\Omega$, and the radial period is $T_{R}=2\pi/\kappa$. (For reference, in a flat rotation curve, we have $\kappa=\sqrt{2}\Omega$ and hence $\gamma=\sqrt{2}$). This epicyclic mapping is accurate up to corrections $\mathcal{O}(\epsilon_{R}^{2})$, where
$\epsilon_{R}(\bm{J})$ $\equiv\left(\frac{J_{R}}{J_{\varphi}}\right)^{1/2}=\frac{1}{\sqrt{\gamma}}\frac{a_{R}}{R_{\mathrm{g}}},$ (51)
is a $\bm{J}$-dependent small quantity, whose typical value is $\sim\epsilon$ (see equation (38)).
It is worth noting the relation between $\langle J_{R}\rangle$ and the mean epicyclic amplitude and radial velocity dispersion. Supposing $f$ is restricted to a narrow enough range of $J_{\varphi}$ values such that $\kappa$ is roughly constant, the root-mean-square epicyclic amplitude is
$a\equiv\sqrt{\langle a_{R}^{2}\rangle}=\left(\frac{2\langle J_{R}\rangle}{\kappa}\right)^{1/2}.$ (52)
Meanwhile, if $f$ is completely phase-mixed (i.e., it has no $\bm{\theta}$-dependence) then we can also straightforwardly relate $\langle J_{R}\rangle$ to the root-mean-square radial velocity:
$\sigma\equiv\sqrt{\langle v_{R}^{2}\rangle}=\left(\kappa\langle J_{R}\rangle\right)^{1/2}.$ (53)
We can also relate $a$ and $\sigma$ to the small parameter (38):
$\epsilon$ $=\sqrt{\langle\epsilon_{R}^{2}\rangle}=\frac{1}{\sqrt{\gamma}}\frac{a}{R_{\mathrm{g}}}=\sqrt{\frac{\gamma}{2}}\frac{\sigma}{V}$ (54)
$\simeq 0.076\times\left(\frac{\gamma}{\sqrt{2}}\right)^{1/2}\left(\frac{\sigma}{20\,\mathrm{km\,s}^{-1}}\right)\left(\frac{V}{220\,\mathrm{km\,s}^{-1}}\right)^{-1},$ (55)
where $V=R_{\mathrm{g}}\Omega$ is the circular speed.
Table 1: Properties of the two orbits shown in Figure 1. Quantities below the single horizontal line are computed using the epicyclic approximation (§3.2).
Orbit (a) (b)
$R_{\mathrm{p}}/R_{0}$ $7.5$ $7.0$
$R_{\mathrm{a}}/R_{0}$ $8.5$ $9.0$
$J_{\varphi}/(R_{0}V_{0})$ $7.97$ $7.90$
$\mathrm{sma}$ $8.0$ $8.0$
$\mathrm{ecc}$ $0.063$ $0.13$
$T_{\varphi}/T_{0}$ $50.1$ $49.6$
$\Omega T_{0}$ $0.13$ $0.13$
$J_{R}/(R_{0}V_{0})$ $0.022$ $0.088$
$a_{R}/R_{0}$ $0.50$ $0.99$
$\epsilon_{R}$ $0.053$ $0.11$
Figure 1: Two orbits in the logarithmic potential. The apo/pericenter distances $(R_{\mathrm{p}},R_{\mathrm{a}})$ for each orbit are shown with red dashed circles, and the initial location is shown with a small blue star. Full details of the orbital parameters are given in Table 1. The top row shows the orbits as viewed in an inertial $(X,Y)$ frame. The bottom row shows the same orbits in the frame rotating azimuthally with frequency $\Omega+J_{R}\partial\,\kappa/\partial J_{\varphi}$ in black. In green we show what the orbits would look like if we ignored the asymmetric drift term and just rotated at $\Omega$.
As an example, we integrated two test-particle orbits in the logarithmic potential
$\widetilde{\phi}(R)=V_{0}^{2}\ln\frac{R}{R_{0}},$ (56)
with $R_{0}$ an arbitrary length scale. This corresponds to a galaxy with a flat rotation curve $V(R)=V_{0}=$ const. (We will employ this potential for examples throughout the paper; to convert to physical units, a sensible choice is $R_{0}=1$ kpc and $V_{0}=220\mbox{ km s}^{-1}$.). In Figure 1 we plot the trajectories of orbits which have the same ‘semimajor axis’ $\mathrm{sma}\equiv(R_{\mathrm{p}}+R_{\mathrm{a}})/2=8R_{0}$, where $(R_{\mathrm{p}},R_{\mathrm{a}})$ are the pericenter/apocenter of the orbit, but different ‘eccentricity’ $\mathrm{ecc}\equiv(R_{\mathrm{a}}-R_{\mathrm{p}})/(R_{\mathrm{a}}+R_{\mathrm{p}})$. Each orbit is integrated for roughly four azimuthal periods, i.e., from time $t=0$ to $t=200\,T_{0}$, where $T_{0}\equiv R_{0}/V_{0}$ is the natural time unit. Further properties of these orbits are summarized in Table 1; in particular, their values of $\epsilon_{R}=(J_{R}/J_{\varphi})^{1/2}$ are $0.053$ and $0.11$ respectively.
The top row of Figure 1 shows these two orbits (a) and (b) as viewed in an inertial frame, while the bottom row shows them in a frame rotating at the frequency $\Omega+J_{R}\,\partial\kappa/\partial J_{\varphi}$, in which the azimuthal frequency is nearly zero [^1] Note this is not quite the same as the azimuthal frequency given in equation (48) because it retains the $\mathcal{O}(\epsilon_{R}^{2})$ asymmetric-drift correction term $J_{R}\,\partial\kappa/\partial J_{\varphi}$. The green lines in Figure 1 show what the orbit would look like if we rotated at frequency $\Omega$, ignoring the drift. While asymmetric drift affects the visual representation of orbits such as those in Figure 1, we find that its dynamical impact is minor, so we mostly ignore it from here on (but see §5.2.2)..
4 Potential fluctuations felt by epicyclic orbits
In this section we introduce the main conceptual advance of our theory, which is to classify potential fluctuations by wavelength (§4.1) and then to find approximate expressions for the Fourier transforms of these fluctuations in each regime (§4.2). We also test our expressions numerically (§4.3).
4.1 Classifying fluctuations by wavelength
The key quantity in our theory is the characteristic radial wavenumber of the $n_{\varphi}$th Fourier component of the guiding center potential perturbation $\delta\Phi_{n_{\varphi}}$:
$k\equiv\frac{1}{\delta\Phi_{n_{\varphi}}}\frac{\partial\delta\Phi_{n_{\varphi}}}{\partial R_{\mathrm{g}}}.$ (57)
Here, and for the remainder of the paper, the $n_{\varphi}$ and $J_{\varphi}$ dependence of this quantity is implicit. Note that $k$ is in general complex (because $\delta\Phi_{n_{\varphi}}$ is in general complex). In particular, in the case that the radial part of the potential takes the form $\delta\Phi_{n_{\varphi}}(R)\propto\exp(ik_{R}R)$ for real $k_{R}$, we have
$k=ik_{R},$ (58)
so that $|k|=|k_{R}|$.
We can now define three regimes of potential perturbations, which will be crucial to the rest of our investigation. Namely, at a fixed $\bm{J}$, we have:
  • Long wavelength regime:
    $|k|\lesssim R_{\mathrm{g}}^{-1},$ (59)
    which is roughly the same as $|k|a_{R}\lesssim\epsilon_{R}$. This regime lies outside the range of validity of the WKB approximation. Typically, only the potentials arising from the lowest-harmonic number, longest-wavelength perturbations like bars, very open spirals, and dwarf galaxies in the outskirts of the disk (like the Large Magellanic Cloud) will fall into this regime.
  • Intermediate wavelength regime:
    $R_{\mathrm{g}}^{-1}\lesssim|k|\lesssim a_{R}^{-1},$ (60)
    or equivalently $\epsilon_{R}\lesssim|k|a_{R}\lesssim 1$. In this case, the potential fluctuations oscillate in space sufficiently rapidly for the WKB approximation to work well, but still slowly enough to be relatively constant on the scale of an individual star’s epicyclic orbit.
  • Short wavelength regime:
    $|k|\gtrsim a_{R}^{-1},$ (61)
    which is approximately equivalent to $|k|R_{\mathrm{g}}\gtrsim\epsilon_{R}^{-1}$. In this regime the WKB approximation works very well, and stars typically feel a potential that oscillates significantly on the scale of their epicyclic orbit.
The way the potential fluctuations couple to stellar orbits will depend strongly upon whether they belong in the long, intermediate or short wavelength regime.
Figure 2: Diagram illustrating the short, intermediate and long wavelength regimes for various epicyclic amplitudes $a_{R}$, or equivalently velocity dispersion of the stellar population $\sigma$ (assuming $a_{R}=a$, see equations (52)-(53)). This plot assumes $R_{\mathrm{g}}=8$kpc, and a flat rotation curve ($\gamma=\sqrt{2}$) with circular velocity $V_{0}=220$ km s-1. We also indicate the typical wavelengths of potential fluctuations due to various astrophysical perturbations.
It is worth writing down an estimate for a typical value of $|k|a$, which equals $\pi$ times the number of radial crests of the potential traversed by a typical star during its epicycle (recall $a$ is the rms value of $a_{R}$, see equation (52)). Using the identities from §3.2, we have:
$|k|a$ $\simeq\sqrt{2}\,\pi\gamma\frac{R_{\mathrm{g}}}{|\lambda|}\frac{\sigma}{V}$ (62)
$\simeq 4.57\times\left(\frac{\gamma}{\sqrt{2}}\right)\left(\frac{R_{\mathrm{g}}}{8\,\mathrm{kpc}}\right)\left(\frac{|\lambda|}{1\,\mathrm{kpc}}\right)^{-1}$
×$\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\times\left(\frac{\sigma}{20\,\mathrm{km\,s}^{-1}}\right)\left(\frac{V}{220\,\mathrm{km\,s}^{-1}}\right)^{-1},$ (63)
where $|\lambda|=2\pi/|k|$.
[…]
Figure 3: Orbit (b) from Figure 1, superimposed on top of the contours of a logarithmic spiral potential (64) with $m=4$ arms and decreasing pitch angles $\alpha$. The pitch angles are chosen such that panels (a), (b) and (c) exhibit potential fluctuations characteristic of the long, intermediate and short wavelength regimes (equations (59)-(61)) respectively.
As a concrete example, in Figure 3 we reproduce the orbit from Figure 1b. In each panel of Figure 3 we superpose this epicyclic orbit on top of a logarithmic spiral potential
$\delta\phi^{\mathrm{ls}}(\varphi,R,t)$
$\equiv-\eta\frac{V_{0}^{2}}{2}\cos\bigg{[}m\cot\alpha\ln\frac{R}{R_{0}}+m(\varphi-\Omega_{\mathrm{p}}t-\varphi_{\mathrm{p}})\bigg{]},$ (64)
which corresponds to a spiral with dimensionless strength $\eta$, $m$ arms (here we set $m=4$), pattern speed $\Omega_{\mathrm{p}}$, pitch angle $\alpha$, and phase $\varphi_{\mathrm{p}}$. For this potential, all power is concentrated at azimuthal wavenumbers $n_{\varphi}=\pm m$. From (57), the corresponding radial wavenumber is:
$k$ $=i\frac{n_{\varphi}\cot\alpha}{R_{\mathrm{g}}}.$ (65)
In particular, $|k|\to 0$ as $\alpha\to 90^{\circ}$ (a very loosely wound spiral), and $|k|\to\infty$ as $\alpha\to 0^{\circ}$ (a very tightly wound spiral). In Figure 3 we illustrate the cases (a) $\alpha=85^{\circ}$,(b) $\alpha=50^{\circ}$ and (c) $\alpha=8^{\circ}$. The resulting values of $|k|a_{R}$ and $|k|R_{\mathrm{g}}$ are given at the top of each panel. Panel (a) is characteristic of the long-wavelength regime (59), panel (b) of the intermediate-wavelength regime (60), and panel (c) of the short-wavelength regime (61).
4.2 Expressions for potential fluctuations in asymptotic regimes
Introducing the three wavelength regimes described above drastically simplifies the approximate expressions for the Fourier components of potential fluctuations $\delta\phi_{\bm{n}}(\bm{J})$. In Appendix B we derive explicit expressions for these functions in each regime up to corrections $\mathcal{O}(\epsilon_{R}^{2})$. Here, we just gather the key results.
Long wavelengths. In this regime the Fourier components generically take the form
$\delta\phi_{\bm{n}}(\bm{J})=\delta\Phi_{n_{\varphi}}(J_{\varphi})\times\begin{cases}1,&n_{R}=0,\\ \pm\frac{n_{\varphi}\gamma a_{R}}{2R_{\mathrm{g}}}\left[1\mp\frac{kR_{\mathrm{g}}}{n_{\varphi}\gamma}\right],&n_{R}=\pm 1,\\ 0,&|n_{R}|\geq 2.\end{cases}$
Thus, for $n_{R}=0$ we simply recover the identity (36); the potential fluctuation acts approximately as if the star were on a perfectly circular orbit. The terms arising from $n_{R}=\pm 1$ are smaller by $\mathcal{O}(\epsilon_{R})$; they are proportional to $J_{R}^{1/2}$, and will ultimately be responsible for Lindblad resonances. Physically, the term involving $n_{\varphi}$ stems from the azimuthal gradients in the potential, while the term proportional to $k$ arises from radial gradients. Finally, for all other $|n_{R}|\geq 2$, long wavelength fluctuations are negligible to the accuracy we are working here.
Short wavelengths. In this regime the potential fluctuation $\delta\phi(\varphi,R)$ can always be expanded as a sum of waves of the form $\psi_{(n_{\varphi},k_{R})}\mathrm{e}^{i(n_{\varphi}\varphi+k_{R}R)}$. The corresponding Fourier components are given in (B7). For tightly wound waves with $|k_{R}|R_{\mathrm{g}}\gg|n_{\varphi}|\gamma$, this reduces to
$\delta\phi_{\bm{n}}(\bm{J})=\psi_{(n_{\varphi},k_{R})}\mathrm{e}^{ik_{R}R_{\mathrm{g}}}(-i\,\mathrm{sgn}\,k_{R})^{n_{R}}J_{n_{R}}\left(|k_{R}|a_{R}\right).$ (67)
However, the tight-winding approximation may not be allowed if one is interested in, e.g., perturbations due to ISM density structures. For instance, in observations of JWST-PHANGS galaxies, Meidt et al. (2023) found much of the power concentrated at $|n_{\varphi}|\sim 50$. For such structures it is better to rely on the more accurate equation (B7).
Intermediate wavelengths. In the intermediate wavelength regime there is technically no uniformly valid asymptotic expansion whose corrections are always $\mathcal{O}(\epsilon_{R}^{2})$. However, if the fluctuation is tightly wound, $|k|R_{\mathrm{g}}\gg|n_{\varphi}|\gamma$, then one can naively extrapolate the result (4.2) from the long-wavelength regime to somewhat shorter wavelengths, or extrapolate the result (67) from the short-wavelength regime to somewhat longer wavelengths. In both cases one arrives at the same answer, namely
$\delta\phi_{\bm{n}}(\bm{J})=\delta\Phi_{n_{\varphi}}(J_{\varphi})\bigg{[}\delta_{n_{R}}^{0}-\frac{1}{2}ka_{R}(\delta_{n_{R}}^{1}+\delta_{n_{R}}^{-1})\bigg{]},$ (68)
where we identify $\delta\Phi_{n_{\varphi}}(J_{\varphi})=\psi_{(n_{\varphi},k_{R})}\mathrm{e}^{ik_{R}R_{\mathrm{g}}}$ (see equation (B6)) and $k=ik_{R}$ (equation (58)). In this sense the intermediate regime provides a smooth interpolation between long and short wavelength regimes, which suggests that — at least for tightly wound waves — the approximate expressions (4.2) and (67) will work well even beyond their strict range of validity.
4.3 Numerical examples
Figure 4: (Real part of the) Fourier components of the potential fluctuations $\delta\phi_{\bm{n}}$ calculated for an $m=4$ logarithmic spiral potential perturbation (64), embedded in a logarithmic background potential, at the action-space location of the orbit shown in Figure 1a (for which $\epsilon_{R}=0.053$). From left to right we decrease the spiral’s pitch angle and hence increase its radial wavenumber $|k|$ (see equation (65)). Panels (a)-(c) show the result for $n_{\varphi}=4$ and for $n_{R}=0$, $1$ and $2$ respectively. The black line shows the result of evaluating the right-hand side of the expression (B2) numerically, while the red and blue dotted lines show the asymptotic results (4.2) and (67) respectively. See text for more details.
Figure 5: As in Figure 4 except evaluated at the action-space location of the orbit shown in Figure 1b. Here $\epsilon_{R}=0.11$ is about twice as large as in Figure 4, leading to a slightly worse agreement between numerical and analytic results.
Figure 6: The solid black, dotted red and dotted blue lines are as in Figure 4 except for $m=25$. The gold dotted line shows the result of using the full WKB expression (B7).
In this subsection we test the accuracy of our asymptotic expressions with some numerical examples.
First, we calculated $\delta\phi_{\bm{n}}(\bm{J})$ for the logarithmic spiral potential (64) with $m=4$ arms, setting $t=0$ and $\varphi_{\mathrm{p}}=0$. To set up the mean field, we used the logarithmic potential (56), and we set $J_{\varphi}=7.97\,R_{0}V_{0}$ and $J_{R}=0.022\,R_{0}V_{0}$, which are the actions of the orbit shown in Figure 1a, and correspond to $a_{R}/R_{0}=0.5$ — see Table 1. We computed $\delta\phi_{\bm{n}}(\bm{J})$ numerically using equation (B2) for different values of the pitch angle $\alpha$, i.e., for different radial wavenumbers $|k|$ (see equation (65)). In Figure 4 we plot with solid black lines the resulting values of Re $\delta\phi_{\bm{n}}(\bm{J})$ (normalized by $\eta V_{0}^{2}/2$) as a function of $|k|$ for $n_{\varphi}=4$ and for $n_{R}=0,1,2$. Note that we have chosen $R_{0}=1$ kpc for this plot, so that wavenumbers are expressed in kpc-1, and orbit in question has a guiding radius of around $8$ kpc and a rms radial velocity of about $20$ km/s.
The formal boundaries between the long, intermediate and short asymptotic wavelength regimes are shown by vertical dashed lines. In the long wavelength regime we overplot the asymptotic result (4.2) with a red dotted line, in the short regime we overplot (67) with a blue dotted line, and in the intermediate regime we overplot both asymptotic results simultaneously. Overall, the agreement between our simple analytic formulae and the exact numerical result is remarkably good, especially at $n_{R}=0$. As $|n_{R}|$ increases, the agreement in the intermediate regime becomes slightly worse, but this is counterbalanced by the fact that intermediate- and long-wavelength fluctuations are suppressed by $\sim|k|\epsilon_{R}^{|n_{R}|}$. Short wavelength fluctuations are not strongly suppressed as $|n_{R}|$ increases, but this causes no trouble since they are very well described by the WKB approximation.
Second, we performed the same calculation at action-space location $J_{\varphi}=7.90\,R_{0}V_{0}$ and $J_{R}=0.088\,R_{0}V_{0}$, which are the actions of the orbit shown in Figure 1b and Figure 3. This orbit has an epicycle size which is twice as large as in Figure 4 (in physical units, it has an rms radial velocity of around $40$ km/s). The agreement between numerical and analytic results, shown in Figure 5, is slightly worse than in Figure 4, which is expected, since $\epsilon_{R}$ has roughly doubled, but is on the whole still excellent.
Third, we experimented with logarithmic spirals of different numbers of arms $m$. We found that the agreement between our simple analytic expressions and the numerical result gets better if we decrease $m$, but gradually gets worse as $m$ is increased. This makes sense mathematically. At short wavelengths a large-$m$ wave is not necessarily tightly wound, so one must replace (67) with the more general form (B7). At long wavelengths, a high-$m$ spiral resembles the spokes of a bicycle wheel: large azimuthal gradients arise even though the radial gradients are small, and so we cannot use the approximation that took us from (B4) to (4.2).
To illustrate these shortcomings we present Figure 6, in which we reproduce the same calculation as in Figure 4 but with $m=25$. We see that the agreement between the blue dotted curves (equation (67)) and the numerical solution is rather poor [^2] The agreement between equation (4.2) (red dotted curves) and the numerics at long and intermediate wavelengths is also poor, but in reality large-$m$ fluctuations nearly always have short wavelengths, so this need not concern us too much.. This can be remedied by using the more accurate non-tightly-wound expression (B7), the result of which is shown with a gold dotted line in Figure 6. In fact, the only difference between (B6) and the full expression (B2) is that we approximate the radial function $\delta\phi_{n_{\varphi}}(R)$ as a sinusoid (equation (B7)). Since logarithmic spirals are locally almost sinusoidal — at each $R$, they take the form $\delta\phi_{n_{\varphi}}(R)\propto\mathrm{e}^{\pm i|k|R\ln R}$ — this approxmation is very accurate. As a result, the gold dashed ‘WKB’ curve reproduces the black numerical curve almost exactly in Figure 6.
Though we have focused on logarithmic spirals for our numerical verification, our results should generalize to other types of perturbations. We conclude that despite their simplicity, the asymptotic results we have derived are remarkably accurate, particularly in the long and short wavelength regimes. Care is needed in the intermediate regime, especially for $n_{R}\neq 0$. Of course, here we have considered only a low-order expansion (ignoring corrections $\mathcal{O}(\epsilon_{R}^{2})$), but more accurate higher order expansions could be developed if so desired.
5 Kinetic theory
Having developed our potential theory in §4, we now illustrate how it can be used to simplify various kinetic calculations. More precisely, we are interested in how the DF $f(\bm{\theta},\bm{J})$ responds to potential perturbations $\delta\phi$ that are assumed ‘small’ in the sense that at all relevant $(\bm{\theta},\bm{J})$,
$\eta\equiv\bigg{|}\frac{\delta\phi}{h_{0}}\bigg{|}\ll 1,$ (69)
where $h_{0}$ is the mean-field Hamiltonian (see equation (14)), with magnitude $|h_{0}|\sim V^{2}/2$).
Linear theory is concerned with changes to the distribution function that are $\mathcal{O}(\eta)$. Next order $\mathcal{O}(\eta^{2})$ effects can often be approached through quasilinear theory, which we discuss in §5.2. However, before we can do any perturbation theory we must acknowledge the fact that we have now introduced two small parameters into our theory, namely $\epsilon\ll 1$ (equation (38)) and $\eta\ll 1$ (equation (69)), but we have not yet described how to order them with respect to each other.
5.1 Ordering small parameters
The parameter $\epsilon$ measures the coldness of the mean-field orbits, while $\eta$ measures the weakness of the fluctuations. If $\epsilon\sim\eta$, we can simply expand our equations (26)-(29) in a single parameter $\epsilon$. The more difficult regimes are (a) $\epsilon\ll\eta$ or (b) $\epsilon\gg\eta$.
First, we deal with regime (a). In this case we argue that fluctuations in the potential will rapidly heat the stellar population until $\epsilon\gtrsim\eta$. To see this, suppose for simplicity that $\epsilon(t=0)=0$, and that $\delta\phi$ is some stochastic (and non-adiabatic) forcing with radial wavelength $\lambda$ and a correlation time $\tau$ (as measured in the frame of a circular orbit). Now ask, what is the value of $\epsilon$ after one orbital period $t_{\mathrm{dyn}}$? There are 3 possible cases to consider depending on the strength and duration of the perturbation:
  • Case I: Suppose the acceleration $\sim|\delta\phi|/\lambda$ is sufficiently strong and/or $\tau$ is sufficiently long that the stars ‘fall down the hill’ of the potential to its bottom. In this case, they pick up a velocity dispersion $\sigma\sim|\delta\phi|^{1/2}$, and hence achieve $\epsilon(t_{\mathrm{dyn}})\sim\sigma/V_{0}\sim\eta^{1/2}\gg\eta$.
  • Case II: If the acceleration is weak and $\tau\sim t_{\mathrm{dyn}}$, then after a correlation time $\tau$ the stars will simply have been kicked by an amount $|\Delta\bm{v}(\tau)|\sim|\delta\phi|\tau/\lambda$. Then the value of $\epsilon$ will have changed by $\Delta\epsilon(\tau)\sim|\Delta\bm{v}(\tau)|/V_{0}\sim\eta V_{0}\tau/\lambda\sim\eta(R_{\mathrm{g}}/\lambda)(\tau/t_{\mathrm{dyn}})$. We can now see that $\epsilon(t_{\mathrm{dyn}})\sim\Delta\epsilon(\tau)\sim\eta(R_{\mathrm{g}}/\lambda)$. Since $R_{\mathrm{g}}/\lambda\gtrsim 1$ for almost all realistic perturbations (see Figure 2), we have $\epsilon(t_{\mathrm{dyn}})\gtrsim\eta$.
  • Case III: If the acceleration is weak and $\tau\ll t_{\mathrm{dyn}}$, following Case II we have that the change in $\epsilon$ after one correlation time is $\Delta\epsilon(\tau)\sim\eta(R_{\mathrm{g}}/\lambda)(\tau/t_{\mathrm{dyn}})$. Then $\epsilon$ will undergo a random walk with step-size $\Delta\epsilon(\tau)$, and so we can estimate $\epsilon(t_{\mathrm{dyn}})\sim\Delta\epsilon(\tau)\times(t_{\mathrm{dyn}}/\tau)^{1/2}\sim\eta(R_{\mathrm{g}}/\lambda)(t_{\mathrm{dyn}}/\tau)^{1/2}$. Using $R_{\mathrm{g}}/\lambda\gtrsim 1$ and $t_{\mathrm{dyn}}/\tau\gg 1$ we again find $\epsilon(t_{\mathrm{dyn}})\gtrsim\eta$.
Second, we consider regime (b), i.e., fluctuations with $\eta\ll\epsilon$. In this case, strictly speaking we ought to extend our epicyclic expansion (§3.2) and the corresponding potential approximations (§4) to higher order in $\epsilon_{R}$ than we have done so far. Suppose for example that $\epsilon\sim\eta^{1/2}$ and we were doing linear theory (i.e., working at $\mathcal{O}(\eta)$). Then we ought to modify both the long (equation (4.2)) and short (equation (67)) wavelength expansions to keep terms $\mathcal{O}(\epsilon_{R}^{2})$. Whether these corrections are important or not depends on the question at hand and the desired accuracy of the prediction. We give an example where $\mathcal{O}(\epsilon_{R}^{2})$ corrections do make a (minor) difference in §5.2.1.
Figure 7: Evolution of $\epsilon\equiv\langle J_{R}/J_{\varphi}\rangle^{1/2}$ for ensembles of $10^{4}$ stars all with the same initial $J_{\varphi}$ but different initial $J_{R}$ when perturbed by a transient spiral (see text for details). The ensembles (shown in different colored curves) are labeled by their initial values of $J_{R}/J_{\varphi}=\epsilon(0)^{2}$, which vary from $10^{-6}$ to $10^{-1}$. Note that the azimuthal period at $J_{\varphi}=8(R_{0}V_{0})$ is $T_{\varphi}\simeq 50T_{0}$). The black dashed line indicates the value $\eta=0.04$; stars with $\epsilon(0)<\eta$ are rapidly heated so that $\epsilon(t_{\mathrm{dyn}})>\eta$.
To test these arguments, we simulated the motion of six ensembles of $10^{4}$ test stars in the logarithmic potential (56), subjected to a transient potential perturbation of the form
$\delta\phi(\varphi,R,t)=\mathrm{e}^{-(t-t_{\mathrm{peak}})^{2}/(2\tau^{2})}\delta\phi^{\mathrm{ls}}(\varphi,R,t),$ (70)
[…]
In Figure 7 we plot the resulting evolution of $\epsilon(t)$ for each of these ensembles, shown with different colored curves, over just the first $\sim 100\,T_{0}$ of the simulation (roughly two azimuthal periods) [^3] Strictly, this measured $\epsilon$ is not exactly equivalent to the $\epsilon$ we defined at fixed $J_{\varphi}$ in equation (38), but we checked that this subtlety has no effect on our conclusions. . To guide the eye, the black dashed line indicates the value of $\eta=0.04$. We see that on this timescale, the $\epsilon$ value for the blue and pink ensembles (those with $\epsilon(0)\gg\eta$, i.e., regime (b)) undergoes only a very modest growth. On the other hand, the red, yellow and green ensembles (those with $\epsilon(0)\ll\eta$, i.e., regime (a)) heat up so that $\epsilon\gtrsim\eta$ within one orbital period. Thereafter, these ensembles exhibit similar modest growth to the initially hotter ensembles.
This behavior turns out to be generic: we found very similar phenomenology when we varied the correlation time $\tau$ and the strength $\eta$ of the perturbation. We have also found the same behavior when integrating test particles in shearing boxes filled with ISM density fluctuations, as we will describe in an upcoming paper.
In summary, either the system is born in the regime $\epsilon\sim\eta$, or (a) it is born with $\epsilon\ll\eta$ but transitions to $\epsilon\sim\eta$ after a time $\lesssim t_{\mathrm{dyn}}$, or (b) it is born with $\epsilon\gg\eta$ and our theory strictly breaks down. For the remainder of this paper we will assume that $\epsilon\sim\eta$ and expand our equations using a single small parameter $\epsilon\ll 1$.
5.2 Quasilinear theory
‘Quasilinear theory’ refers to any kinetic calculation in which an $\mathcal{O}(\eta^{2})$ effect can be computed by multiplying two linear, $\mathcal{O}(\eta)$, fluctuations together. Here we study the quasilinear transport of angular momentum due to a single transient potential perturbation of the form
$\delta\phi(\varphi,R,t)$ $=\,\mathrm{e}^{-(t-t_{\mathrm{peak}})^{2}/(2\tau^{2})}$
×$\,\,\,\,\,\,\,\times A(R)\cos[m\left(\varphi-\Omega_{\mathrm{p}}t+g(R)\right)],$ (71)
for some shape function $mg(R)$, where $m>0$. Throughout this section we ignore any self-gravity of the perturbed part of the DF, i.e., we treat the stars as test particles, so the potential fluctuation is exactly given by (71).
The response of individual stellar orbits to this kind of perturbation has been well-studied (Hamilton & Fouvry, 2024). In particular, a quasilinear description of the orbits is typically valid as long as one can ignore nonlinear trapping at resonances. A rough estimate of the libration period of nonlinearly trapped orbits at corotation can be found using the pendulum approximation (see, e.g., §IV of Hamilton & Fouvry 2024). Namely, perturbations of strength $|A(R_{\mathrm{CR}})/h_{0}|\sim\eta$ (see equation (69)) are trapped at the corotation resonance location $R_{\mathrm{CR}}$ on the timescale
$t_{\mathrm{lib}}$ $\sim 1\,\mathrm{Gyr}\times\bigg{(}\frac{m}{2}\bigg{)}^{-1}\bigg{(}\frac{\eta}{0.01}\bigg{)}^{-1/2}\bigg{(}\frac{t_{\mathrm{dyn}}}{200\,\mathrm{Myr}}\bigg{)}.$ (72)
The pendulum formalism is not really applicable for near-circular orbits at Lindblad resonances, but the libration timescale for these is not too dissimilar from (72).
The calculation we will perform ignores trapping, and so is valid for $\tau\ll t_{\mathrm{lib}}$. It resembles the work of Carlberg & Sellwood (1985), with two small but significant differences. First, our assumption of a Gaussian envelope in (71) is different from their exponential growth and decay, which has the minor advantage that there is no discontinuity in slope at $t=t_{\mathrm{peak}}$. Second, Carlberg & Sellwood (1985) focused on calculating the heating $\Delta\langle J_{R}\rangle$, whereas we are going to focus on the change in the marginalized angular momentum DF $\Delta F_{0}(J_{\varphi})$. An analogous calculation in the opposite limit where $\tau\gg t_{\mathrm{lib}}$ has been performed by Sridhar (2019).
5.2.1 Angular-momentum transport
We first write down a formal expression for the linear response of $f$ to the perturbation $\delta\phi$ by ignoring the right-hand side of (2.2) and integrating in time (assuming an unperturbed initial condition):
$\delta f_{{\bm{n}}}(\bm{J},t)=i{\bm{n}}\cdot\frac{\partial f_{0}}{\partial\bm{J}}\int_{0}^{t}\mathrm{d}t^{\prime}\,\mathrm{e}^{-i{\bm{n}}\cdot\bm{\Omega}(t-t^{\prime})}\delta\phi_{{\bm{n}}}(\bm{J},t^{\prime}).$ (73)
Plugging this in to (32) gives
$\mathcal{F}_{0}$ $=-2\pi\sum_{\bm{n}}n_{\varphi}\int_{0}^{\infty}\mathrm{d}J_{R}\,\bm{n}\cdot\frac{\partial f_{0}}{\partial\bm{J}}$
×$\,\,\,\,\,\,\,\,\,\times\int_{0}^{t}\mathrm{d}t^{\prime}\mathrm{e}^{-i{\bm{n}}\cdot\bm{\Omega}(t-t^{\prime})}\delta\phi_{{\bm{n}}}(\bm{J},t^{\prime})\delta\phi^{*}_{\bm{n}}(\bm{J},t).$ (74)
Let us now specialize to perturbations of the form (71). The Fourier components of this are
$\delta\phi_{\bm{n}}(\bm{J},t)=u_{\bm{n}}(\bm{J})\mathrm{e}^{-in_{\varphi}\Omega_{\mathrm{p}}t}\mathrm{e}^{-(t-t_{\mathrm{peak}})^{2}/(2\tau^{2})},$ (75)
where
$u_{\bm{n}}(\bm{J})=\frac{1}{4\pi}\int{\mathrm{d}\theta_{R}}\mathrm{e}^{-in_{R}\theta_{R}}A(R)\mathrm{e}^{in_{\varphi}[\varphi-\theta_{\varphi}+g(R)]},$ (76)
for $n_{\varphi}=\pm m$, and $u_{\bm{n}}(\bm{J})=0$ otherwise. Plugging (75) into the right-hand side of (74), inserting the resulting expression into (27), and using the epicyclic approximation, we get
$\frac{\partial F_{0}}{\partial t}$ $=2\pi\frac{\partial}{\partial J_{\varphi}}\sum_{n_{\varphi}=\pm m}\sum_{n_{R}}n_{\varphi}\int_{0}^{\infty}\mathrm{d}J_{R}\,|u_{\bm{n}}(\bm{J})|^{2}\,$
×$\,\,\,\,\,\,\,\times\int_{0}^{t}\mathrm{d}t^{\prime}\bm{n}\cdot\frac{\partial f_{0}(\bm{J},t^{\prime})}{\partial\bm{J}}\mathrm{e}^{-i\omega_{\bm{n}}(J_{\varphi})(t-t^{\prime})}\,$
×$\,\,\,\,\,\,\,\times\mathrm{e}^{-(t-t_{\mathrm{peak}})^{2}/(2\tau^{2})}\mathrm{e}^{-(t^{\prime}-t_{\mathrm{peak}})^{2}/(2\tau^{2})},$ (77)
where the frequency
$\omega_{\bm{n}}(J_{\varphi})\equiv n_{\varphi}\Omega+n_{R}\kappa-n_{\varphi}\Omega_{\mathrm{p}},$ (78)
independent of $J_{R}$. Assuming $t_{\mathrm{peak}}\gg\tau$ we can extend the lower limit of the $t^{\prime}$ integration to $\infty$- ∞. Moreover, for sufficiently short-lived perturbations we can ignore the slow time evolution of $f_{0}(\bm{J},t^{\prime})$ on the right-hand side of (77) and integrate directly in time. The result is $F_{0}(J_{\varphi},t\to\infty)=F_{0}(J_{\varphi},0)+\Delta F_{0}(J_{\varphi})$ where
$\Delta F_{0}$ $=\frac{(8\pi^{5})^{1/2}m}{\Gamma}\frac{\partial}{\partial J_{\varphi}}\sum_{n_{R}}\mathcal{R}_{\Gamma}(\omega_{mn_{R}})$
×$\times\int_{0}^{\infty}\mathrm{d}J_{R}\,|u_{mn_{R}}|^{2}\left(m\frac{\partial f_{0}}{\partial J_{\varphi}}+n_{R}\frac{\partial f_{0}}{\partial J_{R}}\right).$ (79)
Here $\Gamma\equiv(\sqrt{2}\tau)^{-1}$ is the width in frequency space associated with the finite lifetime $\tau$ of the perturbation, and
$\mathcal{R}_{\Gamma}(\omega)\equiv\frac{1}{\sqrt{2\pi}\Gamma}\mathrm{e}^{-\omega^{2}/(2\Gamma^{2})},$ (80)
is a Gaussian response function. Finally, one can calculate the change in total angular momentum $J_{\mathrm{tot}}(t)\equiv 2\pi\int\mathrm{d}J_{\varphi}F_{0}(J_{\varphi},t)$ due to the transient by writing
$\Delta J_{\mathrm{tot}}=2\pi\int\mathrm{d}J_{\varphi}\,J_{\varphi}\,\Delta F_{0},$ (81)
and substituting (79) into the right-hand side.
Next we use the results of §4 to evaluate (79) in the long/intermediate wavelength regime, and give a numerical example. We relegate the corresponding short/intermediate wavelength calculation to Appendix C.
5.2.2 Long and intermediate wavelengths
In the long and intermediate wavelength regimes, equations (4.2) and (51) tell us that we may approximate
$|u_{mn}|^{2}=|U_{m}|^{2}\bigg{[}\delta_{n_{R}}^{0}+\frac{J_{R}}{J^{+}_{m}}\delta_{n_{R}}^{1}+\frac{J_{R}}{J^{-}_{m}}\delta_{n_{R}}^{-1}\bigg{]},$ (82)
where $U_{m}(J_{\varphi})\equiv u_{\bm{n}}(J_{\varphi},0)$ (cf. equation (36)) and
$J^{\pm}_{m}(J_{\varphi})\equiv\frac{4J_{\varphi}}{m^{2}\gamma^{3}}\bigg{|}1\mp\frac{kR_{\mathrm{g}}}{m\gamma}\bigg{|}^{-2},$ (83)
is an angular-momentum scale. With this substitution, assuming that $m$ is sufficiently small that $|m\,\partial f_{0}/\partial J_{\varphi}|\ll|\partial f_{0}/\partial J_{R}|$, equation (79) becomes
$\Delta F_{0}=\frac{(8\pi^{5})^{1/2}m}{\Gamma}\frac{\partial}{\partial J_{\varphi}}\left[\Lambda_{m}|U_{m}|^{2}\right],$ (84)
where $\Lambda_{m}=\Lambda_{m}^{0}+\Lambda_{m}^{+}+\Lambda_{m}^{-}$ and
$\Lambda_{m}^{0}=\frac{m}{2\pi}\mathcal{R}_{\Gamma}(\omega_{m,0})\frac{\partial F_{0}}{\partial J_{\varphi}},$ (85)
$\Lambda_{m}^{\pm}=\mp\frac{1}{2\pi}\mathcal{R}_{\Gamma}(\omega_{m,\pm 1})\frac{F_{0}}{J^{\pm}_{m}}.$ (86)
For small $\Gamma$ (large $\tau$) the response function $\mathcal{R}_{\Gamma}$ is sharply peaked around zero frequency. In that case angular-momentum transport is localized near the corotation resonance $\omega_{m,0}=0$, and the inner/outer Lindblad resonances $\omega_{m,\pm 1}=0$. Physically, assuming $\partial F_{0}/\partial J_{\varphi}<0$, stars are transported across the corotation resonance from ‘left to right’, i.e., from $J_{\varphi}<J_{\varphi}^{\mathrm{CR}}$ (where $\omega_{m,0}(J_{\varphi})$ is positive) to $J_{\varphi}>J_{\varphi}^{\mathrm{CR}}$ (where $\omega_{m,0}(J_{\varphi})$ is negative). Overall, the corotation resonance acts to flatten the DF of angular momenta in its own vicinity. On the other hand, the inner (outer) Lindblad resonances always act to push stars to lower (higher) angular momenta, regardless of the gradient of $F_{0}$. This is what allows sharp features like grooves to build up in the angular-momentum DF.
Figure 8: The change in total angular momentum $J_{\mathrm{tot}}$ of a disk of test particles due to an external transient perturbation with characteristic lifetime $\tau$. For details of the setup, see §5.2.2. Solid lines show the results of $10^{6}$-particle simulations. Dashed lines show the analytic results obtained by plugging (84) into (81), while dotted lines correct for asymmetric drift.
[…]
In Figure 8 we show the results of this experiment. With solid curves we plot the simulation results for each decay time $\tau$ in various colors. With dashed lines we show the theoretical prediction for the asymptotic change $\Delta J_{\mathrm{tot}}$ computed by plugging (84) into (81). However, this turns out not to be completely accurate, so with dotted lines we show the corresponding result where we incorporate the effect of asymmetric drift, i.e. we replace $\Omega$ with $\Omega_{\varphi}=\Omega+J_{R}\partial\,\kappa/\partial J_{\varphi}$ in equation (78) [^4] Even though the asymmetric drift correction is $\mathcal{O}(\epsilon^{2})$, the resonance is narrow enough in frequency space (width $\sim\Gamma$) to make it important. Unfortunately, including it means the $J_{R}$ integral must be done numerically.. We see that at the lowest $\tau$, the corrected theory does a very good job of predicting the final $J_{\mathrm{tot}}$ value. As as we increase $\tau$, the theory becomes gradually more inaccurate. At the largest $\tau$ the theory over-predicts the magnitude of $\Delta J_{\mathrm{tot}}$, whereas the numerical result converges to a finite (negative) value. This numerical convergence, and the failure of the theory to predict it, occurs because of the time-dependence of $f_{0}(\bm{J},t)$. This can be induced either by (i) the quasilinear transport, which we ignored in writing down (79), or (ii) orbit trapping at resonances (for the largest $\tau$), which would correspond to a complete breakdown of quasilinear theory. While a detailed study is beyond the scope of this paper, both effects (i) and (ii) produce local flattening of $f_{0}(\bm{J})$ at each resonance, meaning that eventually no more transport can occur there.
6 Conclusions
In this paper we have introduced a scheme for approximating potential fluctuations in galactic disks, and shown how this renders tractable various important calculations from the kinetic theory of stellar systems.
Our findings can be summarized as follows.
  • There are two key lengthscales involved in the description of in-plane orbital motion in a cool galactic disk: the guiding radius $R_{\mathrm{g}}$ and the typical epicyclic amplitude $a$ at that guiding radius. For cool disks, $\epsilon\sim a/R_{\mathrm{g}}\ll 1$.
  • Potential fluctuations in cool disks can be split into long, intermediate and short wavelength regimes relative to $R_{\mathrm{g}}$ and $a$. The Fourier-transformed potential fluctuation $\delta\phi_{\bm{n}}(\bm{J})$, which appears in all sophisticated kinetic theories, and is normally a computational and conceptual bottleneck, is drastically simplified in each asymptotic regime. We gave explicit asymptotic expressions $\delta\phi_{\bm{n}}(\bm{J})$ that are accurate up to relative corrections $\mathcal{O}(\epsilon^{2})$, and we verified these expressions numerically using the example of logarithmic spirals.
  • Our approach also renders tractable various quasilinear expressions that in the past have often been evaluated numerically. As an example, we calculated the quasilinear response of a cool stellar disk to a transient spiral perturbation. We verified the accuracy of this calculation via numerical simulation.
Admittedly, the latter example — the quasilinear response to a transient spiral — was already reasonably well-understood analytically by previous authors (Carlberg & Sellwood, 1985). However, the real power of our galactokinetic scheme is that it will allow us to investigate (semi-)analytically what has previously been studied only by brute-force numerics. In an upcoming paper we will use it to address the crucial issue of radial heating and migration in the Milky Way.
Of course, our theory has its limitations. First, it ignores corrections $\mathcal{O}(\epsilon^{2})$ and higher, which can be on the order of several percent in the Galaxy. Higher order terms are also vital if one wishes to incorporate resonances with $|n_{R}|\geq 2$ into the long/intermediate wavelength theory. One natural extension would therefore be to develop the epicyclic approximation and the associated potential expansion to the next order in $\epsilon$. Second, our quasilinear theory ignores nonlinearly trapped orbits. A next step would be to try to extend the theory to include multiple orbit families, perhaps using the pendulum formalism. Third, we have ignored the self-consistent field $\delta\phi$ produced by the perturbation $\delta f$. It would be desirable to have a self-consistent potential theory derived using the same asymptotic orderings as we used in §4.
Regardless of these shortcomings, if it does indeed prove possible to develop a truly predictive kinetic theory of galactic disks — one that can be meaningfully compared with observations and numerical simulations — then galactokinetics is surely the place to start.
We thank Eve Ostriker for many helpful conversations. C.H. is supported by the John N. Bahcall Fellowship fund at the Institute for Advanced Study. S.M. acknowledges support from the National Science Foundation Graduate Research Fellowship under Grant No. DGE-2039656. This research was supported in part by grant no. NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP).
Appendix A Explicit Fourier expansions
For ease of notation we drop all $\bm{J}$ and $t$ arguments in $\Omega_{\varphi}$, $\delta f$ and $\delta\phi$. Then we have
$\overline{\left(\Omega_{\varphi}\frac{\partial\delta f}{\partial\theta_{\varphi}}\right)}=2\pi\sum_{n_{\varphi}}in_{\varphi}\mathrm{e}^{in_{\varphi}\theta_{\varphi}}\int_{0}^{\infty}\mathrm{d}J_{R}\,\Omega_{\varphi}\,\delta f_{(n_{\varphi},0)},$ (A1)
$\overline{\left(\frac{\partial\delta\phi}{\partial\theta_{\varphi}}\frac{\partial f_{0}}{\partial J_{\varphi}}\right)}=2\pi\sum_{n_{\varphi}}in_{\varphi}\mathrm{e}^{in_{\varphi}\theta_{\varphi}}\int_{0}^{\infty}\mathrm{d}J_{R}\,\frac{\partial f_{0}}{\partial J_{\varphi}}\,\delta\phi_{(n_{\varphi},0)},$ (A2)
$\mathcal{G}=2\pi\sum_{n_{\varphi},n^{\prime}_{\varphi}}\mathrm{e}^{i(n_{\varphi}+n^{\prime}_{\varphi})\theta_{\varphi}}\sum_{n_{R}}\int_{0}^{\infty}\mathrm{d}J_{R}\,\delta f_{(n_{\varphi},n_{R})}\frac{\partial\delta\phi_{(n_{\varphi}^{\prime},-n_{R})}}{\partial J_{\varphi}},$ (A3)
$\mathcal{F}=-2\pi\sum_{n_{\varphi},n^{\prime}_{\varphi}}in^{\prime}_{\varphi}\mathrm{e}^{i(n_{\varphi}+n^{\prime}_{\varphi})\theta_{\varphi}}\sum_{n_{R}}\int_{0}^{\infty}\mathrm{d}J_{R}\,\delta f_{(n_{\varphi},n_{R})}\delta\phi_{(n_{\varphi}^{\prime},-n_{R})}.$ (A4)
Appendix B Fourier components of potential fluctuations in asymptotic regimes
We need the Fourier coefficients $\delta\phi_{\bm{n}}(\bm{J})$ for a given potential $\delta\phi(\varphi,R)$. Inverting (31), the general formula is
$\delta\phi_{\bm{n}}(\bm{J})=\frac{1}{(2\pi)^{2}}\int\mathrm{d}\bm{\theta}\,\delta\phi(\varphi(\bm{\theta},\bm{J}),R(\bm{\theta},\bm{J}))\,\mathrm{e}^{-i\bm{n}\cdot\bm{\theta}}.$ (B1)
Plugging in the epicyclic formulae (42)–(43) we find that, up to corrections $\mathcal{O}(\epsilon_{R}^{2})$,
$\delta\phi_{\bm{n}}(\bm{J})=\sum_{\ell=-\infty}^{\infty}J_{\ell}\left(n_{\varphi}\gamma\frac{a_{R}}{R_{\mathrm{g}}}\right)\int_{0}^{2\pi}\frac{\mathrm{d}\theta_{R}}{2\pi}\mathrm{e}^{-i({n_{R}}-\ell)\theta_{R}}\delta\phi_{{n_{\varphi}}}\left(R_{\mathrm{g}}-a_{R}\cos\theta_{R}\right),$ (B2)
where $J_{\ell}$ is a Bessel function and $\delta\phi_{n_{\varphi}}(R)\equiv(2\pi)^{-1}\int\mathrm{d}\varphi\,\delta\phi(\varphi,R)\mathrm{e}^{-in_{\varphi}\varphi}$ is the $n_{\varphi}$th azimuthal Fourier component of the real-space potential $\delta\phi$ (note this integral is over $\varphi$, not $\theta_{\varphi}$). We get different explicit forms on the right-hand side of this equation depending on which asymptotic wavelength regime we are in.
B.1 Long-wavelength regime
Let us work at a fixed $\bm{J}$ and $\bm{n}$. Supposing that we are in the long wavelength regime (59) we have $|k|a_{R}\lesssim\epsilon_{R}$ (where $|k|$ is defined in (57)). This means that we can expand $\delta\phi_{n_{\varphi}}$ (which was defined below (B2)) in powers of $ka_{R}$ and truncate after the first two terms:
$\delta\phi_{{n_{\varphi}}}(R_{\mathrm{g}}-a_{R}\cos\theta_{R})=\delta\Phi_{n_{\varphi}}(J_{\varphi})(1-{ka_{R}}\cos\theta_{R}+...),$ (B3)
where the discarded terms are $\mathcal{O}\left(|ka_{R}|^{2}\right)\lesssim\mathcal{O}(\epsilon_{R}^{2})$. Then we can plug the expansion (B3) into (B2) and perform the integral over $\theta_{R}$, with the result
$\delta\phi_{\bm{n}}(\bm{J})=\delta\Phi_{n_{\varphi}}(J_{\varphi})\bigg{\{}J_{n_{R}}\left(n_{\varphi}\gamma\frac{a_{R}}{R_{\mathrm{g}}}\right)-\frac{1}{2}ka_{R}\left[J_{n_{R}-1}\left(n_{\varphi}\gamma\frac{a_{R}}{R_{\mathrm{g}}}\right)+J_{n_{R}+1}\left(n_{\varphi}\gamma\frac{a_{R}}{R_{\mathrm{g}}}\right)\right]\bigg{\}}.$ (B4)
Usually in the long wavelength regime, the relevant $n_{\varphi}$ values are of order unity; then since $a_{R}/R_{\rm g}\ll 1$ we can use the small-argument expansion of the Bessel function for $\ell\geq 0$ and $x\geq 0$,
$J_{\ell}(x)\to\frac{1}{\ell!}\left(\frac{x}{2}\right)^{\ell}.$ (B5)
Using this together with the identities $J_{-\ell}(x)=J_{\ell}(-x)=(-1)^{\ell}J_{\ell}(x)$, we find that at long wavelengths the potential fluctuations are given by (4.2).
B.2 Short-wavelength regime
Suppose instead that we are in the short wavelength regime (61), so that that $|k|R_{\mathrm{g}}\gtrsim\epsilon_{R}^{-1}$. Then we are justified in using a WKB approximation, meaning that we may expand an arbitrary potential fluctuation as a sum of sinusoidal waves. In particular we can express the azimuthal Fourier components of the potential fluctuation (defined below (B2)) as
$\delta\phi_{n_{\varphi}}(R)=\psi_{(n_{\varphi},k_{R})}\mathrm{e}^{ik_{R}R},$ (B6)
where $\psi_{(n_{\varphi},k_{R})}$ and $k_{R}$ depend only weakly on $J_{\varphi}$. Corrections to this approximation will be of order $\mathcal{O}\left((|k|R_{\mathrm{g}})^{-2}\right)\lesssim\mathcal{O}(\epsilon_{R}^{2})$, which is negligible to the accuracy we are working with here.
We now insert (B6) into (B2) and use the epicycle approximations (42)–(43). Employing Graf’s addition theorem for Bessel functions (Abramowitz & Stegun, 1968), we find that up to corrections $\mathcal{O}(\epsilon_{R}^{2})$:
$\delta\phi_{\bm{n}}(\bm{J})=\psi_{(n_{\varphi},k_{R})}\mathrm{e}^{i(k_{R}R_{\mathrm{g}}+n_{R}\beta)}(-i\,\mathrm{sgn}\,k_{R})^{n_{R}}J_{n_{R}}\left(K_{R}a_{R}\right),$ (B7)
where we defined the modified wavenumber $K_{R}>0$ and angle $\beta\in(0,\pi/2)$:
$K_{R}\equiv\sqrt{k_{R}^{2}+\frac{n_{\varphi}^{2}\gamma^{2}}{R_{\mathrm{g}}^{2}}},\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\beta\equiv\arctan\left(\frac{|n_{\varphi}|\gamma}{|k_{R}|R_{\mathrm{g}}}\right),$ (B8)
both of which depend on $k_{R}$, $n_{\varphi}$ and $J_{\varphi}$, but not on $J_{R}$.
These expressions simplify further if we may assume that $|n_{\varphi}|$ is sufficiently small so that $|k_{R}|R_{\mathrm{g}}\gg|n_{\varphi}|\gamma$, which is the tightly wound limit. In that case $K_{R}\simeq|k_{R}|$ and $|k_{R}|R_{\mathrm{g}}\gg|n_{R}|\beta$, and we arrive at equation (67).
Appendix C Quasilinear response to a short/intermediate wavelength transient
As a counterpart to §5.2.2, here we consider the angular-momentum transport driven by perturbations of short and intermediate wavelength. It is crucial to consider the short and intermediate wavelength regimes together because, as is clear from Figure 2, a perturbation with, e.g., $|\lambda|\sim 1$ kpc will be of intermediate wavelength from the point of view of the least-eccentric (smallest $a_{R}$) orbits, and of short wavelength from the point of view of more eccentric orbits. The calculation we want to perform, namely evaluating the right-hand side of equation (79), requires us to integrate over all $a_{R}$ values, and so necessarily includes contributions from both regimes.
The key common feature of the short and intermediate wavelength regimes is that $|k|R_{\mathrm{g}}\gtrsim 1$, so we may use a WKB approximation. For simplicity we restrict ourselves to a single value of $k_{R}$, which is equivalent to assuming that $A(R)$ is perfectly sinusoidal, and assume tightly wound waves ($|k_{R}R_{\mathrm{g}}|\gg\gamma m$) so we can use equation (67). Substituting the right-hand side of this equation into (79) and changing integration variable using $J_{R}=\kappa a_{R}^{2}/2$, we get
$\Delta F_{0}(J_{\varphi})=\frac{2\sqrt{2}}{\Gamma}\frac{\partial}{\partial J_{\varphi}}\sum_{n_{R}}\mathcal{R}_{\Gamma}(\omega_{mn_{R}})|\psi_{(n_{\varphi},k_{R})}|^{2}\int_{0}^{\infty}\mathrm{d}a_{R}\,\kappa a_{R}\,\left(m\,\frac{\partial f_{0}}{\partial J_{\varphi}}+n_{R}\,\frac{\partial f_{0}}{\partial J_{R}}\right)|J_{n_{R}}(|k_{R}|a_{R})|^{2}.$ (C1)
To make this integral tractable we suppose the DF has an exponential form in $J_{R}$:
$f_{0}(\bm{J})=\frac{1}{2\pi\langle J_{R}\rangle}F_{0}(J_{\varphi})\,\mathrm{e}^{-J_{R}/\langle J_{R}\rangle}.$ (C2)
Then with the help of the identity $\int_{0}^{\infty}\mathrm{d}x\,x\,\mathrm{e}^{-x^{2}}|J_{\ell}(bx)|^{2}=\frac{1}{2}\mathrm{e}^{-b^{2}/2}\mathcal{I}_{\ell}\left(\frac{b^{2}}{2}\right)$, where $\mathcal{I}_{\ell}(x)$ is the modified Bessel function, we find that $\Delta F_{0}$ is again given by equation (84), except that the function $\Lambda_{m}$ must be modified to
$\Lambda_{m}(J_{\varphi})\equiv\frac{1}{\pi^{5/2}}\mathrm{e}^{-k_{R}^{2}a^{2}/2}\sum_{n_{R}}\mathcal{I}_{n_{R}}\left(\frac{k_{R}^{2}a^{2}}{2}\right)\,\mathcal{R}_{\Gamma}(\omega_{mn_{R}})\left(m\frac{\partial F_{0}}{\partial J_{\varphi}}-n_{R}\frac{F_{0}}{\langle J_{R}\rangle}\right),$ (C3)
where $a^{2}=2\langle J_{R}\rangle/\kappa$ (see equation (52)), and the sum is over all possible $n_{R}$.
The expression (C3) simplifies if we further specify whether the fluctuations are of short wavelength or of intermediate wavelength. However, we must emphasize a subtlety, which is that here we are measuring $k_{R}$ relative to the root-mean-square epicyclic amplitude $a$, i.e., relative to the bulk of the stellar population, rather than to the epicyclic amplitude $a_{R}(\bm{J})$ associated with a particular action $\bm{J}$. If $m$ is large, this sense the names ‘short wavelength’ and ‘intermediate wavelength’ do not have quite the same meanings as we defined back in §4.1.
Consider first the intermediate wavelengths (which now means $|k_{R}|a\ll 1)$. For small arguments $x\ll n_{R}$ the modified Bessel function is well-approximated by $\mathcal{I}_{n_{R}}(x)\to(x/2)^{n_{R}}/n_{R}!$. Thus for $|k_{R}|a\ll 1$ — but still $|k_{R}|R_{\mathrm{g}}\gg 1$ — we have contributions only from $n_{R}=0,\pm 1$. A small amount of algebra will convince the reader that in this limit we in fact recover the same result as in §5.2.2, as expected. Second, consider the short wavelengths. For large arguments $x\gg|n_{R}|$, the modified Bessel function becomes $\mathcal{I}_{n_{R}}(x)\to(2\pi x)^{-1/2}\mathrm{e}^{x}$. This approximation will be inapplicable at sufficiently large $|n_{R}|$, but in those cases the resonance condition is hard to satisfy (assuming $m$ is not large), so the corresponding $\mathcal{R}_{\Gamma}(\omega_{mn_{R}})$ will typically be small. We may therefore ignore large $|n_{R}|$ values and for $|k_{R}|a\gg 1$ find
$\Lambda_{m}(J_{\varphi})\simeq\frac{1}{\pi^{3}|k_{R}|a}\sum_{n_{R}}\mathcal{R}_{\Gamma}(\omega_{mn_{R}}(J_{\varphi}))\left(m\frac{\partial F_{0}}{\partial J_{\varphi}}-n_{R}\frac{F_{0}}{\langle J_{R}\rangle}\right).$ (C4)
We emphasize though that these tightly wound calculations require $m$ not be too large, which is not necessarily true for, e.g., ISM substructure. If $m$ is large, we should not use (67); instead we should use the more general equation (B7), which amounts to replacing the argument of the Bessel function in equation (C1) with $K_{R}a_{R}$ (see the definition (B8)), and hence replacing $k_{R}\to K_{R}$ in (C3).